What are the applications of binomial distribution?
It models the outcome of a number of independent trials in which each trial has only one outcome [that is of interest] with a constant probability of that outcome.
There are random processes that meet these requirements exactly as well as others that may be approximated by the distribution.
What is the probability of getting a 4 or a 6 on one roll of a single die?
2 in 6 or 1 in 3
There is 1 in 6 chance to get a 4 and a 1 in 6 chance to get a 6, combined it is 2 in 6
How many people have the same birthday in the US?
EVERYBODY in the US has the same birthday as someone else in the US. The converse would be that someone in the US has his/her own birthday and nobody else in the entire population of 300,000,000 people in the US was born on that day.
What is the probability that no girls?
We would need to know the number of children in the family to answer this question. For instance, the probability of having no girls in a family of two children would be 1/4 theoretically. In general it is 2-n where n is the number of children.
What does it mean if the mean is doubled but the standard deviation is the same?
What word is used to measure the likelihood that an event will happen?
I think that the word you are looking for is 'probability.'
What are the probable causes for no fuel quantity indication on a Cessna 310 q?
That there is no fuel in it!
What are the formulas for Theoretical Probability?
There are no generic answers. The theoretical probability for rolling a die and tossing a coin will, obviously, be different. The theoretical probability of an event is calculated by finding a suitable model for the trial and then using scientific laws to determine the probabilities of its outcomes.
What is the probability of six?
The answer will depend on what the experiment is: rolling a die, spinning a spinner, the number of times someone will lose before they win (or the converse), the number of rooms in a house, or whatever. Since you have not bothered to share that crucial bit of information, I cannot provide a more useful answer.
What is the probability that on 4 flips of a coin 3 will be heads and1 will be tails?
There are 16 permutations of a coin tossed 4 times. They are ...
HHHH - HHHT - HHTH - HHTT
HTHH - HTHT - HTTH - HTTT
THHH - THHT - THTH - THTT
TTHH - TTHT - TTTH - TTTT
... and the permutations with only one tail are bolded. There are 4, so the probability of flipping a coin 4 times and getting only 1 tail is 4 in 16, or 1 in 4, or 0.25.
The coin does not have a memory therefore the probability is always 50/50.
When two dice are rolled find the probability of getting a 4 or a 10?
When you roll a dice you have these combinations
Die 1:
1 2 3 4 5 6
Die 2:
1 2 3 4 5 6
it can go:
1,1
1,2
1,3
1,4
1,5
1,6
2,1
etc.
which leads to 36 total situations
the only ways to get a 4 are:
1,3
3,1
2,2
2,2
so 4/36
the only ways to get a 10 are:
4,6
6,4
5,5
5,5
4/36
so your chances of 4 or a 10
are 4/36 + 4/36
= 8/36
= 4/18
= 2/9
Drawing a heart from a deck of cards?
There are only 4 suits in an ordinary deck of playing cards. The probability of drawing a heart is 1 in 4, or 0.25.
In a bag of 32 marbles there are 12 blue marbles what proportion of the marbles in the bag is blue?
pedaphile
What questions are asked in the four-way test of the ethical implications of a potential action?
Is the decision truthful?Is it fair to all concerned?Will it build goodwill and better friendships?Will it be beneficial to all concerned?
How many outcomes are possible if you spin the spinner 3 times?
You need to know how many outcomes you have. Is the spinner composed of colors, numbers, names? What categories does the spinner have?
It is 1 - 365Cn/365n.
This is greater than 0.5 for n greater than or equal to 23.
What is the probability of selecting a red hearts or diamonds seven from a deck of card?
It is 14/52 = 7/26.
What is requirement of asymptote in everyday life?
(1)asymptote concept use in spece suttle
(2) landing to perasout
What is the probability of getting a sum of 5 if two dice are tossed 120 times?
Prob(5 in a throw of two dice) = 5/36Prob(Not 5 in a throw of two dice) = 1 - 5/36 = 31/36
Prob(Not 5 in 120 throws) = (31/36)120= 0.000 000 016
and so
Prob(5 in 120 throws) = 1 - Prob(Not 5 in 120 throws) = 0.999 999 984
Prob(5 in a throw of two dice) = 5/36Prob(Not 5 in a throw of two dice) = 1 - 5/36 = 31/36
Prob(Not 5 in 120 throws) = (31/36)120= 0.000 000 016
and so
Prob(5 in 120 throws) = 1 - Prob(Not 5 in 120 throws) = 0.999 999 984
Prob(5 in a throw of two dice) = 5/36Prob(Not 5 in a throw of two dice) = 1 - 5/36 = 31/36
Prob(Not 5 in 120 throws) = (31/36)120= 0.000 000 016
and so
Prob(5 in 120 throws) = 1 - Prob(Not 5 in 120 throws) = 0.999 999 984
Prob(5 in a throw of two dice) = 5/36Prob(Not 5 in a throw of two dice) = 1 - 5/36 = 31/36
Prob(Not 5 in 120 throws) = (31/36)120= 0.000 000 016
and so
Prob(5 in 120 throws) = 1 - Prob(Not 5 in 120 throws) = 0.999 999 984
Can a binomial experiment be used to find the probability of 4 outcomes of .20 and 16?
If the question is about 4 successful outcomes out of 16 trials, when the probability of success in any single trial is 0.20 and independent of the outcomes of other trials, then the answer is, yes, the binomial experiment can be used.
dependent means that a person depends on other person. e.g:-a child depends on his/her parents